OpenStudy (anonymous):

Complete the trinomial so it is a perfect square. 16x2 + ___+ 25

OpenStudy (anonymous):

10x is the missing blank

OpenStudy (anonymous):

40x

OpenStudy (anonymous):

squareroot 25=5*2=10

OpenStudy (anonymous):

How did you get 40x?

OpenStudy (anonymous):

10/2=5 5^2=25

OpenStudy (asnaseer):

shouldn't the answer be 7? 16x2=32 32+25=57 57+7=64 which is a perfect square

OpenStudy (anonymous):

simplify (4x+5)^2

OpenStudy (asnaseer):

is the "x" in this question a variable or the multiplication symbol? I took it to be a multiplication sign.

OpenStudy (anonymous):

a variable

OpenStudy (anonymous):

rld did u get why its 40x?

OpenStudy (anonymous):

ya Thnaks

OpenStudy (asnaseer):

in that case the answer is indeed 40x

OpenStudy (anonymous):

Im confused is it 40 or 10 or what? :S

OpenStudy (anonymous):

y aI totally messed up!!!

OpenStudy (anonymous):

its 40x

OpenStudy (anonymous):

to be precise its plus or minus 40x

OpenStudy (asnaseer):

\[(4x+5)^{2}=16x^{2}+40x+25\]as arjuna indicated.

OpenStudy (anonymous):

or (4x-5)^2

OpenStudy (asnaseer):

true:\[(4x-5)^{2}=16x^{2}-40x+25\]

OpenStudy (asnaseer):

sorry:\[\pm40x\]

OpenStudy (anonymous):

It should be 40x only

OpenStudy (asnaseer):

yes rid613, as the question implies it is "+": 16x2 + ___ + 25